OpenStudy (anonymous):

An object with a charge of +6.0 μC is 0.30 m from a second object and experiences an attractive force of 1.80 N. What is the magnitude of the charge on the second object? A. -6.0 μC B. -3.0 μC C. +3.0 μC D. +6.0 μC

OpenStudy (michele_laino):

here we have to apply the Coulomb law, so we get: \[\large {Q_2} = \frac{{F{d^2}}}{{K{Q_1}}}\]

OpenStudy (anonymous):

ok! so we plug in ?

OpenStudy (michele_laino):

\[\large {Q_2} = \frac{{F{d^2}}}{{K{Q_1}}} = \frac{{18 \times {{0.3}^2}}}{{9 \times {{10}^9} \times 6 \times {{10}^{ - 6}}}}\]

OpenStudy (anonymous):

ok! so we get this? 3E-5?

OpenStudy (michele_laino):

that's right!

OpenStudy (anonymous):

so that would be choice B? :/

OpenStudy (michele_laino):

yes! since the force is attractive

OpenStudy (anonymous):

yay!! thanks!!:D

OpenStudy (michele_laino):

:)